A population of values has a normal distribution with u = 56.1 and o = 57.7. If a random sample of size n = 20 is selected,


- Find the probability that a single randomly selected value is greater than 63.8. Round your answer to four decimals

P(X > 63.8) =


-Find the probability that a sample of size n = 20 is randomly selected with a mean greater than

P(M > 63.8) =





1
Expert's answer
2022-05-05T16:58:42-0400

1)

"P(X>63.8)=1-P(Z\le\dfrac{63.8-56.1}{57.7})"

"\approx1-P(Z\le0.133449)\approx0.4469"

2)

"P(\bar{x}>63.8)=1-P(Z\le\dfrac{63.8-56.1}{57.7/\sqrt{20}})"

"\approx1-P(Z\le0.5968015)\approx0.2753"